A new test was compared with a gold standard measurement with the following results

Sensitivity

Specificity

Positive predictive value

Negative predictive value

New Test

Gold Standard Positive

Gold Standard Negative

Positive

18

2

Negative

8

72
Answer #1

Sensitivity = P(positive on new test and gold standard positive) / P(gold standard positive)

= 18/(18+8)

= 0.6923

= 69.23%

Specificity = P(negative on new test and gold standard negative) / P(gold standard negative)

= 72/(2+72)

= 0.9730

= 97.30%

Positive predictive value = P(positive on new test and gold standard positive) / P(gold standard positive)

= 18/(18+2)

= 0.90

= 90%

Negative predictive value = P(negative on new test and gold standard negative) / P(gold standard negative)

= 72/(8+72)

= 0.9

= 90%